Suppose we make our Physics laboratory a 10 micro-Henry inductor by wrapping a 0.030cm diameter wire around a 4.0 mm diameter cyclical core. How long must this core be?


1
Expert's answer
2021-08-26T14:48:24-0400
"L=\frac{\mu_0 \pi d^2}{4l}\\10^{-5}=\frac{4\pi(10^{-7}) \pi (0.004)^2}{4l}\\l=1.6\cdot10^{-6}\ m"